# Mixed Membership Estimation for Weighted Networks

### v2

@@ 6.4 Real Data
Karate club zachary1977information Member Tie No Known 34 2
Dolphins dolphins0 Dolphin Association No Known 62 2
UKfaculty nepusz2008fuzzy Faculty Friendship No Known 79 3
Political books Krebs (unpublished) Book Co-purchase No Known 92 2
Political blogs Polblogs1 Blog Hyperlink No Known 1222 2
Slovene Parliamentary Party ferligoj1996analysis Party Political space distance Yes Unknown 10 2
Gahuku-Gama subtribes read1954cultures Tribe Friendship Yes Known 16 3
Karate-club-weighted zachary1977information Member Tie Yes Known 34 2
Train bombing hayes2006connecting Terrorist Contact Yes Unknown 64 Unknown
Les Misérables knuth1993stanford Character Co-occurence Yes Unknown 77 Unknown
US Top-500 Airport Network colizza2007reaction Airport #Seats Yes Unknown 500 Unknown
@@
In addition to the computer-generated networks, we also consider some real-world datasets. Table 1 displays basic information for real-world networks considered in this article. For the five un-weighted networks in Table 1, their adjacency matrices and true memberships can be downloaded from http://zke.fas.harvard.edu/software/SCOREplus/Matlab/datasets/. For Slovene Parliamentary Party, it can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/soc/Samo/Stranke94.htm. For Gahuku-Gama subtribes, it can be downloaded from http://konect.cc/networks/ucidata-gama/ and its community information is shown in Figure 9 (b) yang2007community . Karate-club-weighted is the weighted version of Karate club and can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/ucinet/ucidata.htm#kazalo. For Train bombing and Les Misérables, they can be downloaded from http://konect.cc/networks/ (see also kunegis2013konect ). For US Top-500 Airport Network, it can be downloaded from https://toreopsahl.com/datasets/#online_social_network.
@@
For the task of determining the number of communities, we apply KScD, KSPACL, KMixed-SCORE, and their competitors to real-world networks in Table 1. Here, for comparison, we also consider another three approaches BHa, BHm, and BHmc designed based on the Bethe Hessian matrix in le2022estimating . Figure 4 displays the overlapping weighted modularity via Equation (7) using ScD (and SPACL and Mixed-SCORE) for different choices of the number of communities. From Figure 4, we can directly find the $k$ maximizing $Q$ for each method. Table 2 shows the estimated number of communities of KScD, KSPACL, KMixed-SCORE, and their competitors for real networks considered in this paper. For networks with known $K$, KScD, KSPCAL, and KMixed-SCORE correctly infer $K$ for all networks except for the Dophins network while NB, BHac, BHa, BHm, and BHmc fail to determine $K$ for Political blogs and all weighted networks. Again, the fact that our methods significantly outperform NB, BHac, BHa, BHm, and BHmc for the task of inferring $K$ on both un-weighted and weighted real-world networks, suggests the effectiveness and advantage of our overlapping weighted modularity in measuring the quality of overlapping community detection. For Train bombing, Les Misérables, and US Top-500 Airport Network, our methods designed based on overlapping weighted modularity infer their $K$ as 2 while $K$ estimated by NB, BHac, BHa, BHm, and BHmc is larger. As a result, for networks with unknown $K$, we tend to believe that $K$ estimated by our KScD, KSPACL, and KMixed-SCORE is more accurate than their competitors because our approaches outperform their competitors for both simulated weighted networks and real-world networks with known $K$.
@@
Figure 4: Overlapping weighted modularity $Q$ computed by Equation (7) against the number of clusters for real data used in this paper.
For the task of estimating memberships, Table 3 reports the overlapping weighted modularity $Q$ of ScD, GeoNMF, SPACL, and Mixed-SCORE. The results show that all methods have similar or the same $Q$ for all networks. In particular, from Table 3, we find that $Q$ of Karate-club-weighted is larger than that of Karate club, which supports the statement in newman2004analysis that weights provide more useful information about community structure than edges.
@@
Table 3: Overlapping weighted modularity $Q$ (calculated by Equation (7) of ScD and its competitors for real-world networks in Table 1. For networks with known $K$, all four methods use the true $K$ to compute $Q$. For networks with unknown $K$, to compute $Q$, all methods use $K$ estimated by KScD in Table 2. For each data, the largest $Q$ is marked in bold.
@@
  • 10.
    For Les Misérables, it has $77\times 0.0130\approx 1$ highly mixed node and $77\times 0.8052\approx 62$ highly pure nodes. It is the most unbalanced network among all networks because the value 0.3335 of its equilibrium index is much smaller than that of other networks.
@@ 7 Conclusion
Overlapping community detection plays an important role in understanding latent community memberships of complex networks. In this article, we present a general, flexible, and interpretable generative model the Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model for overlapping weighted networks. DCMMDF can be seen as extensions of several previous models in different ways. In particular, the DCMM model is a sub-model of our DCMMDF. The proposed model can model weighted networks in which nodes belong to multiple communities and edge weights can be generated from any distribution as long as the expectation adjacency matrix has a block structure related to the community memberships.
@@
We use an efficient spectral algorithm to estimate the community memberships of weighted networks generated from DCMMDF. We show that the algorithm is asymptotically consistent under mild conditions for any distribution and we can obtain a theoretical guarantee of consistency for a specific distribution from our main theoretical results immediately by analyzing the variance parameter related to the distribution. We present some examples of how to analyze the algorithm’s performance when the adjacency matrix is generated from different distributions under our DCMMDF.
@@
We also present overlapping weighted modularity to evaluate the quality of overlapping community partition for overlapping weighted networks with positive and negative edge weights. We propose a method to determine the number of communities based on the overlapping weighted modularity. Experiments support our theoretical analysis and show that some classical spectral algorithms also successfully estimate community memberships for weighted networks generated from the proposed model. The advantage and effectiveness of our overlapping weighted modularity are supported by the experiments results that our method successfully estimates the number of communities for weighted networks and it outperforms its competitors.
@@
Our work has several implications: First, ground-truth community memberships of overlapping weighted networks generated from DCMMDF provide a reliable way for evaluating mixed membership community detection approaches. Second, algorithms studied in this paper can broaden our understanding of community structure of weighted networks. Third, our overlapping weighted modularity sheds light on evaluating the quality of overlapping community partition for weighted networks with positive and negative edge weights. And last, our KScD provides a baseline for determining the number of communities for weighted networks.
@@
Future works will be studied from four aspects. First, methods with theoretical guarantees should be designed to infer the number of communities for weighted networks generated from DCMMDF. Second, a theoretical guarantee of Mixed-SCORE under DCMMDF should be developed. Third, given a weighted network, analyzing which distribution that edge weights are coming from is an interesting and challenging task. And last, extending DCMMDF and overlapping weighted modularity from un-directed weighted networks to directed weighted networks is appealing.
CRediT authorship contribution statement
The sole author of the paper has performed all the work.
Declaration of competing interest
The author declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Data availability
Data will be made available on request.
@@ Appendix A Proof of Theorem 1
Let $H_{\hat{U}}=\hat{U}^{\prime}U$. Let $H_{\hat{U}}=U_{H_{\hat{U}}}\Sigma_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$ be the top $K$ singular value decomposition of $H_{\hat{U}}$. Let $\mathrm{sgn}(H_{\hat{U}})=U_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$. Under the DCMMDF, the following results are useful for our theoretical analysis.
  • 1.
    $\mathbb{E}[A(i,j)-\Omega(i,j)]=0$ under the DCMMDF.
  • 2.
@@
  • 4.
    Let $\mu=\frac{n\|U\|^{2}_{2\rightarrow\infty}}{K}$ be the incoherence parameter of $\Omega$ defined in Equation (4.23) chen2021spectral . By Lemma H.1 of MaoSVM and Lemma A1 of Criterion , we have $\frac{\theta^{2}_{\mathrm{min}}}{\theta^{2}_{\mathrm{max}}K\lambda_{1}(\Pi^{\prime}\Pi)}\leq\|U\|^{2}_{2\rightarrow\infty}\leq\frac{\theta^{2}_{\mathrm{max}}}{\theta^{2}_{\mathrm{min}}\lambda_{K}(\Pi^{\prime}\Pi)}$, which gives $\frac{\theta^{2}_{\mathrm{min}}n}{\theta^{2}_{\mathrm{max}}K^{2}\lambda_{1}(\Pi^{\prime}\Pi)}\leq\mu\leq\frac{\theta^{2}_{\mathrm{max}}n}{\theta^{2}_{\mathrm{min}}K\lambda_{K}(\Pi^{\prime}\Pi)}$. By Condition 1, we have $\mu=O(1)$.
  • 5.
@@
The above five bullets guarantee that Equations (4.28) and (4.29) of Assumption 4.1. in chen2021spectral hold. Therefore, by Theorem 4.2. chen2021spectral , with probability exceeding $1-O(\frac{1}{n^{5}})$, we have
@@
Since $\sigma_{K}(\Omega)=\sigma_{K}(\Theta\Pi P\Pi^{\prime}\Theta)\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(\Pi P\Pi^{\prime})\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)$ and $\sigma_{1}(\Omega)=\sigma_{1}(\Theta\Pi P\Pi^{\prime}\Theta)\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(\Pi P\Pi^{\prime})\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(P)\lambda_{1}(\Pi^{\prime}\Pi)$ under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, by Condition 1, we have $\kappa(\Omega)=O(1)$, which gives
$\varpi=O(\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\gamma}+\theta_{\mathrm{max}}\sqrt{\gamma\mathrm{log}(n)}}{\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)})=O(\frac{\kappa(\Omega)\sqrt{\gamma}+\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(P)n\sqrt{\rho}}).$ (8)
Corollary 2 of Criterion is a distribution-free theoretical result, it says that there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that for $i\in[n]$,

### v3

@@ 6.4 Real Data
Karate club [59] Member Tie No Known 34 2
Dolphins [82] Dolphin Association No Known 62 2
UKfaculty [77] Faculty Friendship No Known 79 3
Political books Krebs (unpublished) Book Co-purchase No Known 92 2
Political blogs [83] Blog Hyperlink No Known 1222 2
Slovene Parliamentary Party [40] Party Political space distance Yes Unknown 10 2
Gahuku-Gama subtribes [58] Tribe Friendship Yes Known 16 3
Karate-club-weighted [59] Member Tie Yes Known 34 2
Train bombing [84] Terrorist Contact Yes Unknown 64 Unknown
Les Misérables [85] Character Co-occurence Yes Unknown 77 Unknown
US Top-500 Airport Network [86] Airport #Seats Yes Unknown 500 Unknown
CE-CX [87] Gene Line Yes Unknown 15063 Unknown
WormNet-v3 [87] Gene Line Yes Unknown 16258 Unknown
Human-gene1 [88] Gene Interaction Yes Unknown 21853 Unknown
@@
In addition to the computer-generated networks, we also consider some real-world datasets. In this study, we are interested in the largest connected component since real-world networks may have several disconnected components. Table 3 displays basic information for real data considered in this article. For the five un-weighted networks in Table 3, their adjacency matrices and true memberships can be downloaded from http://zke.fas.harvard.edu/software/SCOREplus/Matlab/datasets/. For the Slovene Parliamentary Party network, it can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/soc/Samo/Stranke94.htm. For Gahuku-Gama subtribes, its community information is shown in Figure 9 (b) [89]. Karate-club-weighted is the weighted version of Karate club and can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/ucinet/ucidata.htm#kazalo. For US Top-500 Airport Network, it can be downloaded from https://toreopsahl.com/datasets/#online_social_network. The last three networks in Table 3 are biological networks and they can be downloaded from https://networkrepository.com/bio.php [90]. To simplify our analysis, we treat these real networks as assortative, i.e., $\eta_{\mathcal{N}}=1$ for these networks.
@@
For the task of determining the number of communities, we apply KScD, KSPACL, KMixed-SCORE, and their competitors to real-world networks in Table 3. Here, for comparison, we also consider another three approaches BHa, BHm, and BHmc designed based on the Bethe Hessian matrix in [80]. Figure 8 displays the overlapping weighted modularity via Equation (13) using ScD (and SPACL and Mixed-SCORE) for different choices of the number of communities. From Figure 8, we can directly find the $k$ maximizing $Q_{ovw}$ for each method. Meanwhile, results shown in Table 8 also indicate that the overlapping weighted modularity of ScD is comparable to its competitors as $K$ increases for all data considered in this article. Table 4 shows the estimated number of communities of KScD, KSPACL, KMixed-SCORE, and their competitors for real networks considered in this paper. For networks with known $K$, KScD, KSPCAL, and KMixed-SCORE correctly infer $K$ for all networks except for the Dophins network while NB, BHac, BHa, BHm, and BHmc fail to determine $K$ for Political blogs and all weighted networks. Again, the fact that our methods significantly outperform NB, BHac, BHa, BHm, and BHmc for the task of inferring $K$ on both un-weighted and weighted real-world networks, suggests the effectiveness and advantage of our overlapping weighted modularity in measuring the quality of overlapping community detection. For Train bombing, Les Misérables, and US Top-500 Airport Network, our methods designed based on overlapping weighted modularity infer their $K$ as 2 while $K$ estimated by NB, BHac, BHa, BHm, and BHmc is larger. For CE-CX and Human-gene1, our methods determine their $K$ as 5 and 2, respectively, whereas their competitors fail to output due to memory constraints. For WormNet-v3, KScD and KSPACL determine its $K$ as 7 which is different from that of KMixed-SCORE. As a result, for networks with unknown $K$, we tend to believe that $K$ estimated by our KScD, KSPACL, and KMixed-SCORE is more accurate than their competitors because our approaches outperform their competitors for both simulated weighted networks and real-world networks with known $K$.
@@
(l) CE-CX
(m) WormNet-v3
(n) Human-gene1
Figure 8: Overlapping weighted modularity $Q_{ovw}$ computed by Equation (13) against the number of clusters for real data used in this paper.
For the task of estimating memberships, Table 5 reports the overlapping weighted modularity $Q_{ovw}$ of ScD, GeoNMF, SPACL, and Mixed-SCORE. The results show that all methods have comparable or identical $Q_{ovw}$ values for all networks, except for the CE-CX network. In particular, from Table 5, we find that $Q_{ovw}$ of Karate-club-weighted is larger than that of Karate club, which supports the statement in [39] that weights provide more useful information about community structure than edges. Meanwhile, all methods process the Human-gene1 network of up to 21853 nodes within 14 seconds when there are two communities.
@@
CE-CX Unknown 5 5 5 N/A N/A N/A N/A N/A
WormNet-v3 Unknown 7 7 4 N/A N/A N/A N/A N/A
Human-gene1 Unknown 2 2 2 N/A N/A N/A N/A N/A
@@
CE-CX 0.5120 0.4900 0.4903 0.3566
WormNet-v3 0.2943 0.2552 0.2523 0.2276
Human-gene1 0.3727 0.3754 0.3730 0.3755
Table 5: Overlapping weighted modularity $Q_{ovw}$ (calculated by Equation (13) of ScD and its competitors for real-world networks in Table 3. For networks with known $K$, all four methods use the true $K$ to compute $Q_{ovw}$. For networks with unknown $K$, to compute $Q_{ovw}$, all methods use $K$ estimated by KScD in Table 4. For each data, the largest $Q_{ovw}$ is marked in bold.
@@
  • For Les Misérables, it has $77\times 0.0130\approx 1$ highly mixed node and $77\times 0.8052\approx 62$ highly pure nodes. It is quite unbalanced for its small equilibrium index.
  • For CE-CX and WormNet-v3, both of them have a large proportion of highly mixed nodes and they are even more unbalanced than Les Misérables.
  • For Human-gene1, most of its nodes are highly pure and it is much more balanced than the other two biological networks CE-CX and WormNet-v3.
@@ 7 Conclusion
Overlapping community detection plays an important role in understanding latent community memberships of complex networks. In this article, we present the Weighted Degree-Corrected Mixed Membership (WDCMM) model for overlapping weighted networks. WDCMM releases DCMM’s distribution restriction by allowing the elements of the adjacency matrix to be generated from distributions more than Bernoulli. In this way, WDCMM can model weighted networks in which nodes belong to multiple communities and edge weights can be generated from any distribution as long as the expectation adjacency matrix $\Omega$ has a block structure related to the community memberships.
@@
We use an efficient spectral algorithm to estimate the community memberships of weighted networks generated from WDCMM. We show that the algorithm is asymptotically consistent under mild conditions for any distribution and we can obtain a theoretical guarantee of consistency for a specific distribution from our main theoretical results immediately by analyzing the variance parameter related to the distribution. We present some examples of how to analyze the algorithm’s performance when the adjacency matrix is generated from different distributions under our WDCMM.
@@
We also present overlapping weighted modularity to evaluate the quality of overlapping community partition for both assortative and dis-assortative weighted networks with positive and negative edge weights. We propose a method to determine the number of communities based on the overlapping weighted modularity. Experiments support our theoretical analysis and show that some classical spectral algorithms also successfully estimate community memberships for weighted networks generated from the proposed model. The advantage and effectiveness of our overlapping weighted modularity are supported by substantial experimental results that our method successfully estimates the number of communities for weighted networks and it outperforms its competitors.
@@
Our work has several implications: First, the ground truth community memberships of overlapping weighted networks generated from WDCMM provide a reliable way for evaluating mixed membership community detection approaches. Second, the algorithms studied in this paper can broaden our understanding of the community structure of weighted networks. Third, our overlapping weighted modularity sheds light on evaluating the quality of overlapping community partitions for both assortative and dis-assortative weighted networks with positive and negative edge weights. And last, our KScD provides a baseline for determining the number of communities for weighted networks.
@@
Future works will be studied from five aspects. First,a theoretical guarantee of Mixed-SCORE under WDCMM should be developed. Second, given a weighted network, analyzing which distribution the edge weights are coming from is an interesting and challenging task. Third, extending WDCMM and overlapping weighted modularity from un-directed weighted networks to directed weighted networks is appealing. Fourth, the ScD algorithm and its competitors can be accelerated by some random-projection techniques [91] to detect mixed memberships of large-scale weighted networks. Fifth, extending the higher-order spectral clustering developed in [50] from weighted SBMs to the WDCMM model introduced in this work is appealing. And last, in this paper, we estimate $K$ by maximizing our overlapping weighted modularity for both assortative and dis-assortative networks. Nevertheless, this method fails to infer $K$ in cases where the elements of $P$ are not restricted. Consequently, the development of a more general method, equipped with theoretical guarantees, for estimating $K$ in weighted networks generated from models such as WSBMs and our WDCMM, presents a challenging yet appealing avenue for future research.
[…]
Declarations
Conflict of interest The author declares no conflict of interest.
@@ Appendix A Proof of Theorem 1
Let $H_{\hat{U}}=\hat{U}^{\prime}U$. Let $H_{\hat{U}}=U_{H_{\hat{U}}}\Sigma_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$ be the top $K$ singular value decomposition of $H_{\hat{U}}$. Let $\mathrm{sgn}(H_{\hat{U}})=U_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$. Under the WDCMM, the following results are useful for our theoretical analysis.
  • $\mathbb{E}[A(i,j)-\Omega(i,j)]=0$ under the WDCMM.
@@
  • Let $\mu=\frac{n\|U\|^{2}_{2\rightarrow\infty}}{K}$ be the incoherence parameter of $\Omega$ defined in Equation (4.23) [66]. By Lemma H.1 of [35] and Lemma A1 of [62], we have $\frac{\theta^{2}_{\mathrm{min}}}{\theta^{2}_{\mathrm{max}}K\lambda_{1}(\Pi^{\prime}\Pi)}\leq\|U\|^{2}_{2\rightarrow\infty}\leq\frac{\theta^{2}_{\mathrm{max}}}{\theta^{2}_{\mathrm{min}}\lambda_{K}(\Pi^{\prime}\Pi)}$, which gives $\frac{\theta^{2}_{\mathrm{min}}n}{\theta^{2}_{\mathrm{max}}K^{2}\lambda_{1}(\Pi^{\prime}\Pi)}\leq\mu\leq\frac{\theta^{2}_{\mathrm{max}}n}{\theta^{2}_{\mathrm{min}}K\lambda_{K}(\Pi^{\prime}\Pi)}$. By Condition 1, we have $\mu=O(1)$.
@@
The above five bullets guarantee that Equations (4.28) and (4.29) of Assumption 4.1. in [66] hold. Therefore, by Theorem 4.2.1 [66], with probability exceeding $1-O(\frac{1}{n^{5}})$, we have
@@
Since $\sigma_{K}(\Omega)=\sigma_{K}(\Theta\Pi P\Pi^{\prime}\Theta)\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(\Pi P\Pi^{\prime})\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)$ and $\sigma_{1}(\Omega)=\sigma_{1}(\Theta\Pi P\Pi^{\prime}\Theta)\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(\Pi P\Pi^{\prime})\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(P)\lambda_{1}(\Pi^{\prime}\Pi)$ under $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$, by Condition 1, we have $\kappa(\Omega)=O(1)$, which gives
$\varpi$ $=O(\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\gamma}+\theta_{\mathrm{max}}\sqrt{\gamma\mathrm{log}(n)}}{\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)})=O(\frac{\kappa(\Omega)\sqrt{\gamma}+\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})$
$=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(P)n\sqrt{\rho}}).$ (14)
Corollary 2 of [62] is a distribution-free theoretical result, it says that there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that for $i\in[n]$,